OpenStudy (anonymous):

What is the circumference of a circle with a diameter equaling 7 centimeters in length? Use 3.14 for π

OpenStudy (anonymous):

Circumference = 2(pi)(diameter) C = 2(pi)(7) C = 14pi

jigglypuff314 (jigglypuff314):

no Circumference = 2(radius)(pi) = (diameter)(pi) = (7)(3.14) =

jigglypuff314 (jigglypuff314):

http://www.mathopenref.com/circumference.html